SUB-MILLIMETER WAVE LINES OF THE LOWEST ROTATIONAL TRANSITION OF CH RADICAL

T. AMANO, Institute for Astrophysics and Planetary Sciences, Ibaraki University, Mito, Japan 310-8512.

The lowest rotational transition of CH radical, N = 1, J = 3/2, F_{1} ← N = 1, J = 1/2, F_{2}, has been observed in 532.7 GHz and 536.8 GHz regions with a sub-millimeter wave spectrometer operated with a phase-locked backward-wave oscillator as a radiation source. The CH radical was generated in a DC glow discharge with a mixture of CH_{4}(∼4mTorr) and He(∼240 mTorr) and all the six hyperfine components were recorded, using frequency-discharge double modulation technique^{a}. These transitions had already been observed by using LMR technique^{b} and the zero field transition frequencies were calculated with 1.5 MHz accuracy by Brown and Evenson^{c}. From our new direct observation, more accurate transition frequencies were determined with typically ±30 kHz uncertainty. Astrophysically these accurate frequencies will be of great use for probing hydrocarbon chemistry in warmer and denser regions of molecular clouds.

aT. Amano, paper RD05, 53rd International Symposium on Molecular Spectroscopy, Columbus, Ohio, USA, June, 1998

bJ. T. Hougen et al, J. Mol. Spectrosc.72, 463-483(1978)

cJ. M. Brown and K. M. Evenson, Astrophys. J.268, L51-L56(1983)